OpenStudy (anonymous):

Help?! Will medal and fan............................. ....Find the area of the trapezoid if AB=√12 ,CD=√27 , and AD=2√3 . Use the formula A=1/2h(b1+b2)

OpenStudy (raden):

diagram ?

OpenStudy (anonymous):

sorry forgot it ahah.

OpenStudy (rajee_sam):

AB and CD are b1 and b2, AD is h. substitute in the formula and find your answer

OpenStudy (raden):

area = 1/2 (AB + CD) * AD area = 1/2 (√12 + √27) * 2√3 area = .... hint : √12 = 2√3 √27 = 3√3